OpenStudy (mirra):

if 500 gramsof the substance were present initially and 400 grams are present 50 years later, how many grams will be present after 200 years?

OpenStudy (anonymous):

100 grams disappear every 50 years, so in 200 years 400 more grams will disappear, leaving you with no substance after 200 years.

OpenStudy (texaschic101):

that makes sense :)

OpenStudy (anonymous):

haha

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

mirra you know you can close the question after its been answered so it doesn't show up on the open question list?

OpenStudy (mirra):

Hmm, more explanations pleasee ?

OpenStudy (mirra):

The answer is ?

OpenStudy (anonymous):

the answer is 0 grams

OpenStudy (mirra):

okay, it doesn't help at all. the answer is wrong -,-

OpenStudy (anonymous):

impossible

OpenStudy (anonymous):

@texaschic101 can it be wrong?

OpenStudy (anonymous):

oh that is not even close

OpenStudy (anonymous):

how

OpenStudy (agent0smith):

This is radioactive decay... it's not linear.

OpenStudy (anonymous):

it decreases as follows; every 50 hears you have \(\frac{40}{50}=.8\) of the original amount

OpenStudy (anonymous):

o.

OpenStudy (mirra):

it must have a calculation. its related to exponential function

OpenStudy (agent0smith):

I'd assume it's radioactive decay and follows this formula\[\Large A = A_0 e^{-kt}\] Ao is initial amount, k is rate, t is time

OpenStudy (anonymous):

\[50,50\times .8, 50\times .8^2, 50\times .8^3,...\]

OpenStudy (anonymous):

every 50 years you have 80% of the amount left, there are 4 50 year periods in 200 years compute \[50\times .8^4\] for your answer (use a calculator)

OpenStudy (agent0smith):

500 gramsof the substance were present initially and 400 grams are present 50 years later Ao = 500 when t=50, A=400 \[\Large 400 = 500 e^{-50k}\]solve for k first Then you need to plug t=200 into the formula \[\Large A = 500 e^{-kt} \]

OpenStudy (anonymous):

@agent0smith this is i think too much work, and probably not what the course had in mind (although i could be wrong) it says "related to exponential function" which my guess is \[50\times (.8)^t\] where \(t\) is a multiple of 50, or perhaps if you want times that are not multiples of 50, \[50\times .8^{\frac{t}{50}}\]

OpenStudy (anonymous):

like is said, i could easily be wrong

OpenStudy (agent0smith):

Fair enough, both methods get the same answer, but yours might be quicker

OpenStudy (anonymous):

it would certainly be quicker in this case because \(200=4\times 50\) but your method works as well, although when you truncate the decimal using logs, there will be some error

OpenStudy (agent0smith):

Yeah yours is faster. The error is pretty negligible though (the amounts are the same to 2 d.p. but i used ~6 decimal places for k)